OpenStudy (gelow):

Match each item in Column A to an answer in Column B. 5. (1 pt) Which answer describes the solution to each equation? Column A Column B 1. 5x + 7 = 1/2 (10x +12) 2. 1/9 (18x + 27) = 1+ x -4 - 3x 3. x + 3/4 = 3/4x + 3/4 4. 4 + 3x = 2x - 10 A. one solution B. no solution C. an infinite number of solutions

OpenStudy (gelow):

@bibby

OpenStudy (gelow):

@KamiBug

OpenStudy (gelow):

@SolomonZelman

OpenStudy (gelow):

@GretaKnows

OpenStudy (gelow):

@pandagirl811

OpenStudy (gelow):

@lizz123

OpenStudy (gelow):



OpenStudy (gelow):

@shelbydancer

OpenStudy (gelow):

@misty1212

OpenStudy (gelow):

@lol23

OpenStudy (gelow):

@KendrickLamar2014

OpenStudy (anonymous):

don't know sorry

OpenStudy (gelow):

@MrBraveDragon

OpenStudy (gelow):

@sammixboo

OpenStudy (anonymous):

Ok

OpenStudy (gelow):

so what do i do?????????????

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Ok im back

OpenStudy (gelow):

@Mendicant_Bias

OpenStudy (gelow):

@Agl202

OpenStudy (anonymous):

Do you do k12?

OpenStudy (gelow):

yes

OpenStudy (anonymous):

What unit?

OpenStudy (anonymous):

are you on?

OpenStudy (gelow):

Unit 6 | Lesson 9

OpenStudy (anonymous):

Just passed it XD

OpenStudy (anonymous):

Ok i will help you

OpenStudy (anonymous):

Which question?

OpenStudy (anonymous):

@gelow Are you there?

OpenStudy (gelow):

5 sorry

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

thats the one i got wrong sorry

OpenStudy (gelow):

it dont telll you the answer?

OpenStudy (anonymous):

nope it's my 4th time taking it xD

OpenStudy (anonymous):

it's a hard test

OpenStudy (anonymous):

atleast you can get a %80

OpenStudy (anonymous):

tell me what you got for 1.

OpenStudy (gelow):

ohh

OpenStudy (anonymous):

What did you get for number 1?

OpenStudy (gelow):

Which step should be completed first to solve the equation ? A. Add 14 to both sides. B. Divide both sides by 3. C. Divide both sides by −3. D. Subtract 9 from both sides.

OpenStudy (gelow):

d

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

What was the equation?

OpenStudy (gelow):

did i get it? oh sorry

OpenStudy (gelow):

9-3q=14

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Number 2?

OpenStudy (gelow):

What is the solution to the equation ? c =

OpenStudy (anonymous):

what is the equation?

OpenStudy (gelow):

3/4 (4c + 16 )= 2c + 9 =-3

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Number 3?

OpenStudy (gelow):

What is the solution to the equation ? m =

OpenStudy (gelow):

-5 8m+12/4 -5

OpenStudy (anonymous):

Im checking 1 sec

OpenStudy (anonymous):

What did you get for m?

OpenStudy (gelow):

ok

OpenStudy (anonymous):

What did you get for m? sorry

OpenStudy (gelow):

-5

OpenStudy (anonymous):

so the equation is -5 8m+12/4?

OpenStudy (anonymous):

I don't think thats -5 is the answer

OpenStudy (anonymous):

I think you put the equation wrong

OpenStudy (anonymous):

Are you there?

OpenStudy (gelow):

yes sorry and why?

OpenStudy (anonymous):

Do you see where it says draw?

OpenStudy (anonymous):

draw me the equation

OpenStudy (gelow):

ok

OpenStudy (gelow):